The dimensions of a cuboid ate acm, acm,2acm. Write an expression for the surface area in terms of a. The surface area is 1000cm. Calculate the value of a. Calculate the volume of the cuboid
Answer by Boreal(15235)

There are six faces, two of which are a x a so that is 2a^2
The other four are 2a x a, or a total of 8a^2
The total surface area is 10a^2
The volume is a x a x 2a=2a^3
10a^2=1000
a^2=100
a=10 ANSWER
volume is 2a^3=2000 cm^3.